OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Slope = 4, passing through (-3, 7)

OpenStudy (anonymous):

I need help..

OpenStudy (lucaz):

point slope form is (y-y1)=m(x-x1) you the slope 4, the point (-3,7), plug them in the equation

OpenStudy (anonymous):

so what's the answer

OpenStudy (anonymous):

y + 7 = 4(x - 3)

OpenStudy (lucaz):

(y-(7))=4(x-(-3)) y-7=4(x+3)

OpenStudy (anonymous):

so which one is the answer @lucaz

OpenStudy (lucaz):

y-7=4(x+3)

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Slope = , 6/7 passing through (8, 7) @lucaz

OpenStudy (anonymous):

I got y - 7 =6/7 (x - 8) as the answer

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 7 = 6/7(x - 8) you got it correct

OpenStudy (lucaz):

now you know how to do it

OpenStudy (anonymous):

This is the answer to your problem and I wrote it all out so you can see how I got the answer.